OpenStudy (anonymous):

F(x)=2+x^2 and g(x)=sqrt(x-1) find f(g(x))

OpenStudy (jdoe0001):

\(\bf f(x)=2+x^2\\ g(x)=\color{blue}{\sqrt{x-1}}\\ f(g(x)) = 2+\left(\color{blue}{\sqrt{x-1}}\right)^2 \)

OpenStudy (anonymous):

real help is appreciated.

OpenStudy (loser66):

@ehamer2990 what does it mean? does it mean jdoe0001's answer helpless?

OpenStudy (anonymous):

to me yes, it is helpless

OpenStudy (loser66):

@jdoe0001 come here to fix it, friend. the asker doesn't want the answer, he wants to know how to solve.

OpenStudy (jdoe0001):

well, so.... \(\bf f(g(x)) = 2+\left(\color{blue}{\sqrt{x-1}}\right)^2\)

OpenStudy (jdoe0001):

anything confusing about it?

OpenStudy (jdoe0001):

\(\bf \left(\sqrt{x-1}\right)^2 \implies \sqrt{(x-1)^2}\)

OpenStudy (anonymous):

why on earth are colors involved!?

OpenStudy (jdoe0001):

am I wrong in assuming you're using Internet Explorer? or some older browser?

OpenStudy (loser66):

lol

OpenStudy (anonymous):

Im using chrome and I even tried firefox, your formatting was wrong, therefore it didnt show up correctly.

OpenStudy (jdoe0001):

hmmm? i shows up in firefox and chrome just fine :/